A wire of 0 m length carries the current I1 = 25 A. Another wire of...

  1. A wire of 0 m length carries the current I1 = 25 A. Another wire of the same length, parallel to the first one carries the current I2 = 15 A. As a result, the wires attract each other with the force F = 5 mN. What is the distance between the wires? Are the currents of the same or opposite directions?

What of the following statements is/are correct?

  1. The current in one wire create electric field, which excerts a force on the second wire with the current;
  2. The current in one wire create magnetic field, which excerts a force on the second wire with the current;
  3. The force between the wires is inverse-proportional to the distance between wires;
  4. The force between the wires is inverse-proportional to the squared distance between wires;
  5. A and D;
  6. B and C.
Answer #1

Force = 2*k*i1*i2*L/R

R = 2*10^-7*25*15*L/5*10^-3

R = 150*L*10^-4 m

Value of L is not given so please put value of L

The force is attractive means current is in same direction in both the wires.

Option F is correct means b and c statement is true.